# Magnetism in Twisted Bilayer WSe2

### v1

@@ Abstract
Conclusions— Our self-consistent Hartree-Fock calculations offer a useful semi-quantitative guide of the ferromagnetic behavior observed experimentally. Our results emphasize the importance of first having the interaction-renormalized band structure before incorporating the effects of strong correlations. In such moiré systems, the interaction driven modification of the band structures can be significant, even at the mean-field level. As illustrated here for the case of tWSe2, experimental observations of correlated states can be understood as a simple Stoner-like ferromagnetism once the interaction-modification of the bands are accounted for. These results highlight the importance of long-range Coulomb interactions in capturing the quantum phases in twisted 2D materials and the methods we employ are generally applicable to a wide class of moiré systems. Several future directions can be pursued within this framework. For instance, the superconducting phase at half-filling could be explored by incorporating superconducting diagrams within the Nambu formalism [41]. Additionally, the reported generalized Wigner crystals at fractional fillings [6] can be investigated by considering other translational symmetry-broken ground states.

### v2

@@ Abstract
Conclusions— Our self-consistent Hartree-Fock calculations offer a useful semi-quantitative guide of the ferromagnetic behavior observed experimentally. Our results emphasize the importance of first having the interaction-renormalized band structure before incorporating the effects of strong correlations. In such moiré systems, the interaction driven modification of the band structures can be significant, even at the mean-field level. As illustrated here for the case of tWSe2, experimental observations of correlated states can be understood as a simple Stoner-like ferromagnetism once the interaction-modification of the bands are accounted for. These results highlight the importance of long-range Coulomb interactions in capturing the quantum phases in twisted 2D materials and the methods we employ are generally applicable to a wide class of moiré systems. Several future directions can be pursued within this framework. For instance, the superconducting phase at half-filling could be explored by incorporating superconducting diagrams within the Nambu formalism [49].